the measure of the supplement of an angle is 12 degrees greater than three times the measure of a complement. Find the measure of the angle.
180-12 = 3(x) but this does not work. the answer is 51 degrees. I don't know how to get it.

Answer by mananth(16946)

the measure of the supplement of an angle is 12 degrees greater than three times the measure of a complement. Find the measure of the angle.
..
let angle be x
complement = 90-x
supplement = 180-x
..
180-x=3(90-x)+12
180-x =270-3x+12
add +x
180-x+x=270-3x+x+12
180=-2x+12 +270
add-12
180-12=-2x+12-12
-102=-2x
x=-102/-2
x=51